ann is 5 years older than mary. the sum of their ages is 35 find ann's age and mary's ages
Answer by T-Maree(17)

Mary: 15 years old
Ann: 20 years old
the difference in age = d = 5 years
35 - d = 30
30 / 2 = 15
therefore mary is 15
15 + d = 20
therefore ann is 20
20 + 15 = 35